OpenStudy (anonymous):

Your family moved to a tropical climate. The first month, you recorded the number of rainy days that occurred each week. Your data contained: 6, 7, 5, 6. Which statement about this sample is true? A) Mode < Mean B) Median > Mode C) Mode > Mean D) Mean = Mode

OpenStudy (texaschic101):

mean is the average and mode is the number that appears most often. To find the mean (average), you add all the numbers together, then divide by how many numbers there are. So the mean is : (6 + 7 + 5 + 6) / 4 = 24/4 = 6 and the mode (the number that appears the most) is 6 Therefore, the mean = the mode